OpenStudy (anonymous):

A hairdresser is deciding where to open her own studio. If the hairdresser chooses Location A, she will pay $1200 per month in rent and will charge $45 per haircut. If she chooses Location B, she will pay $1800 per month in rent and will charge $60 per haircut. How many haircuts would she have to give in one month to make the same profit at either location?

OpenStudy (perl):

need nhelp?

OpenStudy (perl):

A = 1200 + 45x B = 1800 + 60x solve A = B